OpenStudy (anonymous):

8 1/5 - 3 4/10 show yourwork

OpenStudy (anonymous):

\[8 • \frac{ 1}{ 5 } - 3 • \frac{ 4 }{ 10 } = \frac{ 2 }{ 5 }\] How? Because when you simplify 8-3 Which equals 5 then simplifying the numerator and denominator it will leave you with 2/5 (Hope this helped)

OpenStudy (anonymous):

thank you very much it did help